OpenStudy (anonymous):

Multiply and simplify. y^3*y^-5*y^-2*y

OpenStudy (anonymous):

\[3-5-2+1\] will be your exponent

OpenStudy (anonymous):

y^-3

OpenStudy (anonymous):

well the options are - A. y^-4 B. 1/y^4 C. y^-5 D. 1/y^3

OpenStudy (anonymous):

1/y^3

OpenStudy (agreene):

or if you prefer: \[y^3y^{-5}y^{-2}y=\frac{y^3y}{y^5y^2}\] \[\frac{y^3y}{y^5y^2}= \frac{y^4}{y^7}\implies\frac{\cancel{y^4}}{\cancel{y^4}y^3}\] \[\frac{1}{y^3} =y^{-3}\]